It took Steve 6 hours to drive to his house from his cousins at a velocity of 90 KM/H. If he wanted to take 8 hours to make the return trip home, how fast should he drive?
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
d = 90 * 6 = 540 km
s = 540 / 8
Answer:
s = 67.5 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php